# Forecasting causal dynamics with universal reservoirs

### v1

@@ 3.1 The iterated multistep forecasting scheme
Assume also that the readout $h:\mathbb{R}^{L}\longrightarrow{\mathbb{R}}^{d}$ is Lipschitz with constant $L_{h}$. Suppose that $H^{F}_{h}$ is a uniform $\varepsilon$-approximant of $H$ and hence it satisfies (3.1). Let ${\bf y}=({\bf y}_{t})\in({\mathbb{R}}^{d})^{\mathbb{Z}}$ be a realization of the chain $H$, that is, ${\bf y}_{t}=H\left(\underline{{\bf y}_{t-1}}\right)$ for all $t\in\mathbb{Z}$, and let $\widetilde{{\bf y}}_{t}=H^{F}_{h}(\underline{{\bf y}_{0}}\,\underline{\widetilde{{\bf y}}_{t-1}})$, $t\in\mathbb{N}$, be the forecast of the positive terms of ${\bf y}$ using $(F,h)$ and the iterative scheme introduced in (3.2) and (3.3). Let $\widetilde{\bf x}_{t}=H^{F}(\underline{{\bf y}_{0}}\,\underline{\widetilde{{\bf y}}_{t-1}})$ be the corresponding sequence in the state space produced by these recursions and let ${\bf x}_{t}=H^{F}(\underline{{\bf y}_{t-1}})$, a sequence in the state space parallel to ${\bf y}$. Then for any $\Phi$-invariant Borel probability measure on $X^{\prime}$, there is a set $\Lambda\subset X^{\prime}$ of full measure such that, if ${\bf x}_{0}\in\Lambda$, then for sufficiently small $\delta>0$, there exists a constant $R=R(\delta,{\bf x}_{0})$ and a time horizon $T=T(\delta,{\bf x}_{0})\in\mathbb{N}$ such that
@@
Notwithstanding what has been said above, in the error bounds (3.14), (3.15), we do see an interplay between overfitting and approximation error. The top Lyapunov exponent is related to the regularity of the dynamical system generated by the state-space system, while the approximation error $\varepsilon$ determines the set of state-space systems we can choose from. Thus decreasing the approximation error may lead to a highly irregular dynamical system and hence a large leading Lyapunov exponent. See [Berr 23, Remark 12] for further discussion on this.
@@
In this setup, it is visible how there are three contributions to the error bound in (3.15), namely the approximation error $\varepsilon$ of $H$ by $H^{F}_{h}$, the Lipschitz constants $L_{h}$ and $L_{z}$, which are related to the structure of the RNN, and the exponential factor $e^{t(\lambda_{1}^{\text{pos}}+\delta)}$ which is exclusively related to the dynamical features of the map $\Phi\in\operatorname{Diff}^{1}(X)$ that $(F,h)$ determines. Much work has been done on the ability of state-space systems to reproduce the Lyapunov exponents of the dynamical system being learnt. Two factors come into play here: first, while it is known that under suitable conditions on the GS of a state-space system that embeds the dynamical system in a higher dimensional space, the Lyapunov spectrum of the original system is contained in that of the reconstructed system, a number of additional Lyapunov exponents arise, commonly referred to as spurious Lyapunov exponents, some of which may exceed the top Lyapunov exponent of the original dynamical system [Dech 96, Dech 00]; second is the question of continuity of Lyapunov exponents under perturbations of the generating dynamical system, which is discussed in detail in the survey [Vian 20]. Of particular interest, then, is the stability gap, namely the difference between the top Lyapunov exponent of the reconstructed system and that of the original system. An upper bound for the stability gap is given in [Berr 23, Theorem 3 (iii)]. In our case, however, we are concerned only with the Lyapunov exponents of the dynamical system associated to $(F,h)$. Indeed, it is more useful for us to have bounds for the error in terms of the Lyapunov exponents of the learnt dynamical system, which can be calculated, rather than that of the dynamical system we are learning, which are unknown to us. Nevertheless, work on the stability gap is relevant: as discussed in [Berr 23, Theorem 3], the stability gap is always nonnegative, so when learning the observations from a dynamical system, we cannot hope to get a reconstructed dynamical system with lower Lyapunov exponent than that of the original. We note that, in general, CCIMs need not have Lyapunov exponents.
@@
$\left\|\Delta{{\bf y}}_{t}\right\|\leq\varepsilon\left(1+L_{h}L_{z}\sum_{j=0}^{t-2}\left(L_{x}+L_{h}L_{z}\right)^{j}\right),\quad\mbox{for all $t\in\mathbb{N}$.}$ (3.17)
@@
$\left\|\Delta{{\bf y}}_{t}\right\|\leq\varepsilon\left(\frac{1-L_{x}}{1-\left(L_{x}+L_{h}L_{z}\right)}\right),\quad\mbox{for all $t\in\mathbb{N}$.}$ (3.18)
@@
The approximation bound $\varepsilon>0$ in (3.14) and (3.15) which has been chosen so that $\left\|H-H^{F}_{h}\right\|_{\infty}<\varepsilon$ can in practice be put in relation, using Barron-type bounds, with architecture parameters of the RNN families that are being used in the forecasting exercise. Relations of this type can be found in [Gono 23a, Gono 23b].
Proof of the theorem. As in the proof of Theorem 3.1, we take $\Lambda$ to be the set of (Lyapunov-Perron) regular points in $X^{\prime}$ of the dynamical system $\Phi$. Let ${\bf x}_{0}\in\Lambda.$ As noted before, according to the iterations (3.2) and (3.3), the sequence $(\widetilde{\bf x}_{t})$ is a trajectory of the dynamical system $\Phi$, that is, $\widetilde{\bf x}_{t}=\Phi(\widetilde{\bf x}_{t-1})$ for $t\in\mathbb{N}$ with initial condition $\widetilde{\bf x}_{0}={\bf x}_{0}.$ Defining ${\bf C}_{t}:=F({\bf x}_{t-1},{\bf y}_{t-1})-F({\bf x}_{t-1},h({\bf x}_{t-1}))$ for $t\in\mathbb{N},$ we have
@@
$={\bf C}_{t}+\Phi({\bf x}_{t-1})-\Phi(\widetilde{\bf x}_{t-1})={\bf C}_{t}+D\Phi(\widetilde{\bf x}_{t-1})\Delta{\bf x}_{t-1}+{\bf O}_{t},$ (3.20)
@@
$\Delta{\bf x}_{t}={\bf C}_{t}+\sum_{s=1}^{t-1}D\Phi(\widetilde{\bf x}_{t-1})\cdots D\Phi(\widetilde{\bf x}_{s}){\bf C}_{s}+\left({\bf O}_{t}+\sum_{s=2}^{t-1}D\Phi(\widetilde{\bf x}_{t-1})\dots D\Phi(\widetilde{\bf x}_{s}){\bf O}_{s}\right)\\ ={\bf C}_{t}+\sum_{s=1}^{t-1}D\Phi^{t-s}(\widetilde{\bf x}_{s}){\bf C}_{s}+\left({\bf O}_{t}+\sum_{s=2}^{t-1}D\Phi^{t-s}(\widetilde{\bf x}_{s}){\bf O}_{s}\right)={\bf a}_{t}+{\bf O}_{t}+\sum_{s=2}^{t-1}D\Phi^{t-s}(\widetilde{\bf x}_{s}){\bf O}_{s},$ (3.21)
@@
$\leq e^{t(\lambda_{1}^{\text{pos}}+\delta)}\left(\varepsilon L_{z}R+L_{\Phi}K\sum_{s=2}^{t}e^{-s(\lambda_{1}^{\text{pos}}-3\delta)}\lVert\Delta{\bf x}_{s-1}\rVert^{2}\right).$
@@
$b_{t}=e^{t(\lambda_{1}^{\text{pos}}+\delta)}\left(\varepsilon L_{z}R+L_{\Phi}K\sum_{s=2}^{t}e^{-s(\lambda_{1}^{\text{pos}}-3\delta)}b_{s-1}^{2}\right),\quad t\in\mathbb{N};\quad b_{0}=0.$
@@
We can bound the sequence ${\bf c}$ by an exponential growth factor for a finite time horizon. Consider equation (3.22). Suppose that for $s=1,\dots,t-1$ we have $c_{s}\leq 1+\eta,$ for some $\eta>0.$ Then
$c_{t}$ $\leq 1+\varepsilon L_{z}L_{\Phi}KRe^{-2(\lambda_{1}^{\text{pos}}+\delta)}\sum_{s=2}^{t}e^{s(\lambda_{1}^{\text{pos}}+5\delta)}(\eta+1)^{2}\leq 1+(\eta+1)^{2}\varepsilon L_{z}L_{\Phi}KR\left(\frac{e^{8\delta}}{e^{\lambda_{1}^{\text{pos}}+5\delta}-1}\right)e^{(t-1)(\lambda_{1}^{\text{pos}}+5\delta)}.$
@@ 3.2 The causal embedding forecasting scheme
$=f_{2}^{\prime}\circ\phi^{\prime}\circ\phi^{\prime t-1}(m^{\prime})=\left(f^{\prime}(\phi^{\prime t}(m^{\prime})),f^{\prime}(\phi^{\prime t+1}(m^{\prime}))\right)=\left(U^{F^{\prime}}({\bf m}^{\prime})_{t},U^{F^{\prime}}({\bf m}^{\prime})_{t+1}\right).$
@@ 4 Numerical illustration
In this section we illustrate the error bounds given in Theorem 3.11 for predicting observations coming from dynamical systems using reservoir systems. Our results show the applicability of the error bound to these systems in accurately predicting the rate of growth of the forecasting errors. Nevertheless, there are some limitations to the error bound, which become apparent in our numerical work. First of all, the dynamical systems we learn all have attractors with finite diameter, and hence, while the echo state network (ESN) that we use for the forecasting (see 3.19) typically reconstructs the attractor, the actual predictions deviate from the true values over time until the error saturates the attractor, being bounded by its diameter, and hence it settles at a constant value. After this, the error bound is no longer applicable. Second, the error bound given in (3.15) includes the multiplicative constant $R$. The existence of this constant is guaranteed by the use of arguments involving Oseledets’ Theorem in the proofs; intuitively, the constant gives the system time to settle until the ergodic properties underlying this theorem take effect. However, we do not have an algorithm for calculating this constant. Instead, we estimate it by considering the log prediction errors and fitting the intercept of the log error bound to this. In the case of reservoir systems that do not learn the given system well, this constant may turn out to exceed the error at saturation, and the bound becomes redundant. While on first impression this may seem perplexing, this is a practical difficulty inherent to any learning paradigm. Rather than undermining the usefulness of the error bound given in this paper, it serves to show the inability of the given reservoir computer to track the dynamical system, in which case any error bounds are redundant as saturation is quickly reached. We emphasize that in the case of reservoir systems that demonstrate the ability to produce accurate long-term iterated multistep predictions, the gradient of the log error growth closely matches the predicted value given in (3.15). Nevertheless, the question naturally arises, in the context where we have to deal with observations from an unknown dynamical system, whether we can determine up to what time the error bounds are valid. Once again, as noted in Remark 3.14, while we are guaranteed the existence of a finite time horizon $T$ up to which the bound holds, in practice, the horizon is difficult to calculate. In the empirical exercises below, the bound is valid until saturation. If the reservoir system reconstructs the attractor of the dynamical system well, the saturation error can be estimated by running predictions on any two distinct trajectories of the dynamical system with close initial conditions. The time horizon up to which error estimates are both valid and useful can be calculated by checking when the predicted error bound crosses the saturation error. However, this still requires knowledge of the multiplicative constant $R$. Thus, to get a more clear answer to this question, more work needs to be done on calculating this constant.
@@
with an affine readout given by $h\colon\mathbb{R}^{L}\longrightarrow\mathbb{R}^{d},h({\bf x})=W{\bf x}+{\bf a}$ for some linear map $W\in\mathbb{M}_{d,L}$ and bias ${\bf a}\in\mathbb{R}^{d}.$ We take the nonlinear function $\boldsymbol{\sigma}(\cdot)$ to be tanh()($\cdot$)( ⋅ ) applied componentwise and randomly generate the connectivity matrix $A\in\mathbb{M}_{L,L}$ and input mask $C\in\mathbb{M}_{L,d}$ by selecting entries independently from a uniform distribution over the interval $[-1,1]$, with density $0.01$ and $1$, respectively. The input shift $\boldsymbol{\zeta}\in\mathbb{R}^{L}$ is simply taken to be zero. Data for trajectories of the different systems is generated by integrating the given dynamical system over a discrete time step $\Delta t$ using the Runge-Kutta45 method. In each case, an input trajectory ${\bf y}_{-T},\dots,{\bf y}_{0}$ is fed into the randomly generated state map, generating a trajectory in the state space according to the scheme
${\bf x}_{t+1}=F({\bf x}_{t},{\bf y}_{t}),\quad\mbox{for}\quad t=-T,\dots,-1,\quad\mbox{where}\quad{\bf x}_{-T}={\bf 0}.$ (4.2)
An initial washout of $\tau$ timesteps allows the state space trajectory to converge to the values given by the GS so that the two become indistinguishable, although the state space trajectory generated is arbitrarily initialized at $\bf 0$. This is due to the so-called input forgetting property [Grig 19]. The washout data is discarded and a ridge regression with regularization constant $\alpha$ is performed between the input points ${\bf y}_{-T+\tau},\dots,{\bf y}_{0}$ and state space points ${\bf x}_{-T+\tau},\dots,{\bf x}_{0},$ to fit the output parameters $W$ and ${\bf a}$. Predictions are then made by iterating (3.2) and (3.3), that is,
@@
for $t=1,2,3,\dots,$ where $\widetilde{\bf x}_{0}={\bf x}_{0}$ and $\widetilde{\bf y}_{0}={\bf y}_{0}.$ There are several hyperparameters that need to be tuned, namely the regularization constant $\alpha$, the scaling constant $\gamma$ for the input mask, and the spectral radius of the connectivity matrix. Selecting these is a perplexing task as, to our knowledge, no clear analytical method exists by which prediction error as a function of these constants may be minimized. Moreover, the error in prediction is not stable with respect to choices of the constants and, especially in selecting the regularization constant, small changes may lead to largely different results. In order to tune our ESNs individually to each system, we had to resort to a brute force cross validation method minimizing mean squared prediction error over a grid for these paremeters. This code is computationally expensive.

### v2

@@ 3.1 The iterated multistep forecasting scheme
Assume also that the readout $h:\mathbb{R}^{L}\longrightarrow{\mathbb{R}}^{d}$ is Lipschitz with constant $L_{h}$. Suppose that $H^{F}_{h}$ is a uniform $\varepsilon$-approximant of $H$ and hence it satisfies (3.1). Let ${\bf y}=({\bf y}_{t})\in({\mathbb{R}}^{d})^{\mathbb{Z}}$ be a realization of the chain $H$, that is, ${\bf y}_{t}=H\left(\underline{{\bf y}_{t-1}}\right)$ for all $t\in\mathbb{Z}$, and let $\widetilde{{\bf y}}_{t}=H^{F}_{h}(\underline{{\bf y}_{0}}\,\underline{\widetilde{{\bf y}}_{t-1}})$, $t\in\mathbb{N}$, be the forecast of the positive terms of ${\bf y}$ using $(F,h)$ and the iterative scheme introduced in (3.2) and (3.3). Let $\widetilde{\bf x}_{t}=H^{F}(\underline{{\bf y}_{0}}\,\underline{\widetilde{{\bf y}}_{t-1}})$ be the corresponding sequence in the state space produced by these recursions and let ${\bf x}_{t}=H^{F}(\underline{{\bf y}_{t-1}})$, a sequence in the state space parallel to ${\bf y}$. Then for any $\Phi$-invariant Borel probability measure on $X^{\prime}$, there is a set $\Lambda\subset X^{\prime}$ of full measure such that, if ${\bf x}_{0}\in\Lambda$, then for sufficiently small $\delta>0$, there exists a constant $R=R(\delta,{\bf x}_{0})>0$ and a time horizon $T=T(\delta,{\bf x}_{0})\in\mathbb{N}$ such that
@@
Another limitation of the bounds in (3.11) and (3.12) is that, while the theorem guarantees the existence of the constant $R=R({\bf x}_{0},\delta)$, it does not tell how to calculate it. We propose a numerical procedure for estimating $R$ in Section 4.
@@
Notwithstanding what has been said above, in the error bounds (3.11)-(3.12), we do see an interplay between overfitting and approximation error. The top Lyapunov exponent is related to the regularity of the dynamical system generated by the state-space system, while the approximation error $\varepsilon$ determines the set of state-space systems we can choose from. Thus decreasing the approximation error may lead to a highly irregular dynamical system and hence a large leading Lyapunov exponent. See [Berr 23, Remark 12] for further discussion on this.
Remark 3.17
In this setup, it is visible how there are three contributions to the error bound in (3.12), namely the approximation error $\varepsilon$ of $H$ by $H^{F}_{h}$, the Lipschitz constants $L_{h}$ and $L_{z}$, which are related to the structure of the RNN, and the exponential factor $e^{t(\lambda_{1}^{\text{pos}}+\delta)}$ which is exclusively related to the dynamical features of the map $\Phi\in\operatorname{Diff}^{1}(X)$ that $(F,h)$ determines. Much work has been done on the ability of state-space systems to reproduce the Lyapunov exponents of the dynamical system being learnt. Two factors come into play here: first, while it is known that under suitable conditions on the GS of a state-space system that embeds the dynamical system in a higher dimensional space, the Lyapunov spectrum of the original system is contained in that of the reconstructed system, a number of additional Lyapunov exponents arise, commonly referred to as spurious Lyapunov exponents, some of which may exceed the top Lyapunov exponent of the original dynamical system [Dech 96, Dech 00]; second is the question of continuity of Lyapunov exponents under perturbations of the generating dynamical system, which is discussed in detail in the survey [Vian 20]. Of particular interest, then, is the stability gap, namely the difference between the top Lyapunov exponent of the reconstructed system and that of the original system. An upper bound for the stability gap is given in [Berr 23, Theorem 3 (iii)]. In our case, however, we are concerned only with the Lyapunov exponents of the dynamical system associated to $(F,h)$. Indeed, it is more useful for us to have bounds for the error in terms of the Lyapunov exponents of the learnt dynamical system, which can be calculated, rather than that of the dynamical system we are learning, which are unknown to us. Nevertheless, work on the stability gap is relevant: as discussed in [Berr 23, Theorem 3], for a state-space system that reconstructs the original dynamical system, the stability gap is always nonnegative. Thus, when learning the observations from a dynamical system, if our trained is conjugate to the underlying dynamical system, we cannot hope to get a reconstructed dynamical system with lower Lyapunov exponent than that of the original. Naturally, due to errors in learning, the trained state-space system will not perfectly reconstruct the dynamical system and we may have a negative stability gap, as illustrated in our numerical section where we train an ESN on the Lorenz system and obtain a maximum Lyapunov exponent lower than that of the true system. In this case, however, we would expect an increase in the other constants involved in the bound, such as $\varepsilon$ and $R$, since if our proxy system has vastly different Lyapunov exponents from that of the true underlying system, it cannot yield a good approximation of it. We note that, in general, CCIMs need not have Lyapunov exponents.
@@
$\left\|\Delta{{\bf y}}_{t}\right\|\leq\varepsilon\bigg{(}1+L_{h}L_{z}\sum_{j=0}^{t-2}\left(L_{x}+L_{h}L_{z}\right)^{j}\bigg{)},\quad\mbox{for all $t\in\mathbb{N}$.}$ (3.13)
@@
$\left\|\Delta{{\bf y}}_{t}\right\|\leq\varepsilon\left(\frac{1-L_{x}}{1-\left(L_{x}+L_{h}L_{z}\right)}\right),\quad\mbox{for all $t\in\mathbb{N}$.}$
@@
The approximation bound $\varepsilon>0$ in (3.11) and (3.12) which has been chosen so that $\left\|H-H^{F}_{h}\right\|_{\infty}<\varepsilon$ can in practice be put in relation, using Barron-type bounds, with architecture parameters of the RNN families that are being used in the forecasting exercise. Relations of this type can be found in [Gono 23, Gono 24].
Proof of the theorem. As in the proof of Theorem 3.1, we take $\Lambda$ to be the set of (Lyapunov-Perron) regular points in $X^{\prime}$ of the dynamical system $\Phi$. Let ${\bf x}_{0}\in\Lambda$. As noted before, according to the iterations (3.2)-(3.3), the sequence $(\widetilde{\bf x}_{t})$ is a trajectory of the dynamical system $\Phi$, that is, $\widetilde{\bf x}_{t}=\Phi(\widetilde{\bf x}_{t-1})$ for $t\in\mathbb{N}$ with initial condition $\widetilde{\bf x}_{0}={\bf x}_{0}$. Defining ${\bf C}_{t}:=F({\bf x}_{t-1},{\bf y}_{t-1})-F({\bf x}_{t-1},h({\bf x}_{t-1}))$ for $t\in\mathbb{N}$, we have
@@
$={\bf C}_{t}+\Phi({\bf x}_{t-1})-\Phi(\widetilde{\bf x}_{t-1})={\bf C}_{t}+D\Phi(\widetilde{\bf x}_{t-1})\Delta{\bf x}_{t-1}+{\bf O}_{t},$
@@
$\Delta{\bf x}_{t}={\bf C}_{t}+\sum_{s=1}^{t-1}D\Phi(\widetilde{\bf x}_{t-1})\cdots D\Phi(\widetilde{\bf x}_{s}){\bf C}_{s}+\bigg{(}{\bf O}_{t}+\sum_{s=2}^{t-1}D\Phi(\widetilde{\bf x}_{t-1})\dots D\Phi(\widetilde{\bf x}_{s}){\bf O}_{s}\bigg{)}\\ ={\bf C}_{t}+\sum_{s=1}^{t-1}D\Phi^{t-s}(\widetilde{\bf x}_{s}){\bf C}_{s}+\bigg{(}{\bf O}_{t}+\sum_{s=2}^{t-1}D\Phi^{t-s}(\widetilde{\bf x}_{s}){\bf O}_{s}\bigg{)}={\bf a}_{t}+{\bf O}_{t}+\sum_{s=2}^{t-1}D\Phi^{t-s}(\widetilde{\bf x}_{s}){\bf O}_{s},$ (3.15)
@@
$\leq e^{t(\lambda_{1}^{\text{pos}}+\delta)}\bigg{(}\varepsilon L_{z}R+L_{\Phi}K\sum_{s=2}^{t}e^{-s(\lambda_{1}^{\text{pos}}-3\delta)}\lVert\Delta{\bf x}_{s-1}\rVert^{2}\bigg{)}.$
@@
$b_{t}=e^{t(\lambda_{1}^{\text{pos}}+\delta)}\bigg{(}\varepsilon L_{z}R+L_{\Phi}K\sum_{s=2}^{t}e^{-s(\lambda_{1}^{\text{pos}}-3\delta)}b_{s-1}^{2}\bigg{)},\quad t\in\mathbb{N};\quad b_{0}=0.$
@@
We can bound the sequence ${\bf c}$ by an exponential growth factor for a finite time horizon. Consider (3.16) and suppose that for $s=1,\dots,t-1$ we have $c_{s}\leq 1+\eta$, for some $\eta>0$. Then
$c_{t}$ $\leq 1+\varepsilon L_{z}L_{\Phi}KRe^{-2(\lambda_{1}^{\text{pos}}+\delta)}\sum_{s=2}^{t}e^{s(\lambda_{1}^{\text{pos}}+5\delta)}(\eta+1)^{2}$
$\leq 1+(\eta+1)^{2}\varepsilon L_{z}L_{\Phi}KR\left(\frac{e^{8\delta}}{e^{\lambda_{1}^{\text{pos}}+5\delta}-1}\right)e^{(t-1)(\lambda_{1}^{\text{pos}}+5\delta)}.$
@@ 3.2 The causal embedding forecasting scheme
$=f_{2}^{\prime}\circ\phi^{\prime}\circ\phi^{\prime t-1}(m^{\prime})=\left(f^{\prime}(\phi^{\prime t}(m^{\prime})),f^{\prime}(\phi^{\prime t+1}(m^{\prime}))\right)$
$=(U^{F^{\prime}}({\bf m}^{\prime})_{t},U^{F^{\prime}}({\bf m}^{\prime})_{t+1}).$
@@ 4 Numerical illustration
In this section we illustrate the error bounds given in Theorem 3.11 for predicting observations coming from dynamical systems using reservoir systems. Our results show the applicability of the error bound to these systems in accurately predicting the rate of growth of the forecasting errors. Nevertheless, there are some limitations to the error bound, which become apparent in our numerical work. First of all, the dynamical systems we learn all have attractors with finite diameter, and hence, while the echo state network (ESN) that we use for the forecasting in (3.14) typically reconstructs the attractor, the actual predictions deviate from the true values over time until the error saturates the attractor, being bounded by its diameter, and hence it settles at a constant value. This is a limitation inherent to the system. After saturation, any error bound becomes redundant.
Second is the question which naturally arises from the first limitation, namely: in the context where we deal with observations from an unknown dynamical system, are we able to determine up to what time the error bounds are both valid and meaningful? Theorem 3.11 guarantees validity of the bound up to a finite time horizon $T$. However, as noted in Remark 3.14, this horizon is difficult to calculate in practice. Notwithstanding this, in the empirical exercises below, it is found that the bound remains valid over the entire forecasted trajectory. More importantly, the fitted bounds in our experiments give us meaningful information about the forecasting error up to saturation. We may calculate the saturation error by running predictions on any two distinct trajectories of the dynamical system with close initial conditions until their separation distance stabilizes. The time horizon up to which error estimates are both valid and useful can then be calculated by checking when the predicted error bound crosses this saturation error.
This brings us to the third though standard limitation, which has to do with the constant $R$ in (3.12), which multiplies the exponential growth term. This factor needs to be calculated in order to fit the error bounds for a given initial condition. The existence of this constant is guaranteed by the use of arguments involving Oseledets’ Theorem in the proofs; intuitively, the constant gives the system time to settle until the ergodic properties underlying this theorem take effect. While the theory does not supply an algorithm for calculating $R$, we instead estimate the constant by selecting it as the minimum amongst all possible values of $R$ for which the bound holds. In view of Theorem 3.11, this may be seen as a minimum estimate for the constant $R$ guaranteed in the theorem. While the constant guaranteed in the theory may be higher, as is often the case in applications, in practice the bound holds for a much lower value of $R$. In the case of reservoir systems that do not learn the given system well, the constant from the theorem may indeed turn out to exceed the error at saturation, and the bound becomes redundant. While on first impression this seems perplexing, this is a practical difficulty inherent to any learning paradigm. Rather than undermining the usefulness of the error bound given in this paper, it serves to demonstrate the inability of the given reservoir computer to track the dynamical system, in which case any error bounds are redundant as saturation is quickly reached.
Despite all these limitations, we emphasize that the usefulness of the bound lies in the fact that the exponential growth rate which ultimately dominates the growth of the error can be calculated, as it depends on the top Lyapunov exponent of the trained system, to which we have access. Indeed, as the numerical experiments below illustrate, in the case of reservoir systems that demonstrate the ability to produce accurate long-term iterated multistep predictions, the gradient of the log error growth of forecasting errors closely matches the predicted value given in (3.12).
@@
with an affine readout given by $h\colon\mathbb{R}^{L}\longrightarrow\mathbb{R}^{d},h({\bf x})=W{\bf x}+{\bf a}$ for some linear map $W\in\mathbb{M}_{d,L}$ and bias ${\bf a}\in\mathbb{R}^{d}$. We take the nonlinear function $\boldsymbol{\sigma}(\cdot)$ to be tanh()($\cdot$)( ⋅ ) applied componentwise and randomly generate the connectivity matrix $A\in\mathbb{M}_{L,L}$ and input mask $C\in\mathbb{M}_{L,d}$ by selecting entries independently from a uniform distribution over the interval $[-1,1]$, with density $0.01$ and $1$, respectively. The input shift $\boldsymbol{\zeta}\in\mathbb{R}^{L}$ is simply taken to be zero. Data for trajectories of the different systems is generated by integrating the given dynamical system over a discrete time step $\Delta t$ using the Runge-Kutta45 method. In each case, an input trajectory ${\bf y}_{-T_{0}},\dots,{\bf y}_{0}$ is fed into the randomly generated state map, generating a trajectory in the state space according to the scheme
${\bf x}_{t+1}=F({\bf x}_{t},{\bf y}_{t}),\quad\mbox{for}\quad t=-T_{0},\dots,-1,\quad\mbox{where}\quad{\bf x}_{-T_{0}}={\bf 0}.$ (4.2)
An initial washout of $\tau$ timesteps allows the state space trajectory to converge to the values given by the GS so that the two become indistinguishable, although the state space trajectory generated is arbitrarily initialized at $\bf 0$. This is due to the so-called input forgetting property [Grig 19]. The washout data is discarded and a ridge regression with regularization constant $\alpha$ is performed between the input points ${\bf y}_{-T_{0}+\tau},\dots,{\bf y}_{0}$ and state space points ${\bf x}_{-T_{0}+\tau},\dots,{\bf x}_{0}$, to fit the output parameters $W$ and ${\bf a}$. Predictions are then made by iterating (3.2) and (3.3), that is,
@@
for $t=1,2,3,\dots,T_{1}$ where $\widetilde{\bf x}_{0}={\bf x}_{0}$ and $\widetilde{\bf y}_{0}={\bf y}_{0}$. There are several hyperparameters that need to be tuned, namely the regularization constant $\alpha$, the scaling constant $\gamma$ for the input mask, and the spectral radius of the connectivity matrix. Selecting these is a perplexing task as, to our knowledge, no clear analytical method exists by which prediction error as a function of these constants may be minimized. Moreover, the error in prediction is not stable with respect to choices of the constants and, especially in selecting the regularization constant, small changes may lead to largely different results. In order to tune our ESNs individually to each system, we had to resort to a brute force cross validation method minimizing mean squared prediction error over a grid for these parameters. This code is computationally expensive.
Computation of constant $R$.
In our empirical exercises we propose to calculate the constant $R$ for the error bound (3.12) as follows: for a given trajectory
$({\bf y}_{-T_{0}},{\bf y}_{-T_{0}+1}\dots,{\bf y}_{-1},{\bf y}_{0},{\bf y}_{1},{\bf y}_{2},\dots,{\bf y}_{T_{1}}),$
after the ESN had been trained and predictions $(\widetilde{\bf y}_{1},\widetilde{\bf y}_{2},\dots,\widetilde{\bf y}_{T_{1}})$ made, $R$ was selected as
$R:=\arg\min_{R^{\prime}>0}\{R^{\prime}\colon\varepsilon\left(1+2L_{h}L_{z}R^{\prime}e^{t\Delta t(\lambda_{1}^{\text{pos}}+\delta)}\right)\geq\|{\bf y}_{t}-\widetilde{\bf y}_{t}\|\quad\text{for all }t=1,\dots,T_{1}\}.$ (4.5)
[…]
From this observation, we also propose a method for calculating $R$ without making use of the testing data: Typically in the learning task at hand one is given data from the trajectories of the system being learnt, which we divide between trajectories for validation and trajectories for testing. Hyperparameters are selected by a cross validation procedure on the validation data. In this procedure trajectories are divided into a training segment $\{{\bf y}_{-T_{0}},{\bf y}_{-T_{0}+1},\dots,{\bf y}_{0}\}$ and a validation segment $\{{\bf y}_{1},{\bf y}_{2},\dots,{\bf y}_{T_{1}}\}$. The ESN is initialized with different hyperparameters, its readout trained on the training portion of the trajectories and the ESN then evaluated by its performance in predicting the rest of the trajectories. Finally the best hyperparameter combination is chosen. Following this, the testing trajectories are divided in the same way as the validation trajectories, the second segment this time being for testing. The ESN readout is trained on the training portion of the data and predictions made over the testing segment. To calculate a value of $R$ which would give a reasonable bound for most trajectoires, we may run the ESN with trained readout on the validation data, calculate $R$ for each trajectory using (4.5), and then take an average over these values. As seen in the figures below, with initial conditions coming from the same distribution as those used during validation, the bound accounts for approximately $95\%$ of the trajectories from this distribution.
The main emphasis of the figures remains to illustrate how the gradient of the log error growth matches that given by the bound, namely the top Lyapunov exponent of the trained system.